A school festival has sold 760-tickets. If this is 95% of their goal, how many more tickets do they need to sell to reach 100% of their goal?
A. 22 B. 40 C. 58 D. 800
: A school festival has sold 760-tickets. If this is 95% of their goal, how many more tickets do they need to sell to reach 100% of their goal?
A. 22 B. 40 C. 58 D. 800

Answer by mangopeeler07(445)

760/x=95/100-----------------------760 is 95% of what number?

Cross multiply:
76000=95x

Divide by 95:
800=x

So the goal is 800. You already sold 760, so you need
800-760
more tickets

800-760=40
40 more tickets

Answer:
b)40